OpenStudy (anonymous):

solve and show work 4(c-3)-6=10 c=?

OpenStudy (anonymous):

distribute the 4 within the paranthesis to get 4c-12-6=10. Then combine your like terms to get c.

OpenStudy (anonymous):

you will get 4c = 28 then just divide each term by 4 and c = 7

OpenStudy (anonymous):

Expand: 4(c)=4c 4(-12)-6=10 4c-18=10 add 18 to both sides 4c=28 divide each side by 4 c=7

OpenStudy (anonymous):

now thts more lik it